OpenStudy (hpfan101):

\[\lim_{x \rightarrow 0} \frac{ \sin5x }{ 3x }\]

OpenStudy (anonymous):

Hint:\[\frac{\sin5x}{3x}=\frac{5}{3}\times\frac{\sin5x}{5x}\]

OpenStudy (hpfan101):

Ok, thank you! I was able to figure it out! :)